The DaqBoxes 11 and 09 (SDB2_LeftDown and SDB2_LeftUp) as well as the DaqBox 58 (SQZ_DBOX_2) were no more sending their data packets since 13h22m25 UTC.
We suspect that there has been an issue in the timing 100 MHz distribution to the DET elec lab, to be confirmed with further analysis.
We first reconfigured the SQZ_DBOX_2. Then we stopped/start the power supply of the LeftDown and LeftUp daqboxes and could recover them. Since the timing inside SDB2 is propagated in series started from Dbox11, we have reconfigured the timing of all the DaqBoxes. At this occasion, we found that there has been a shift of 10 ns since 6th November on all DaqBoxes on 100 MHz (so it was transparent for demodulation phase): we have reconfigured all the DaqBoxes (SDB2, LNFS, SSFS, SIB2, SPRB) so that they are all in sync again.